OpenStudy (anonymous):

Divide sqrt 180x^3/sqrt 10x^-7

OpenStudy (anonymous):

\[\LARGE \frac{\sqrt{180x^3}}{\sqrt{10x^{-7 }}}\] hint... split 180 into prime numbers...

OpenStudy (anonymous):

I got 3sqrt2/2x^2 which was incoorect

OpenStudy (anonymous):

\[\LARGE \frac{\sqrt{180}\cdot \sqrt{x^2}}{\sqrt{10}\cdot \sqrt{x^{-7} }}=\] \[\LARGE \frac{\sqrt{90\cdot 2 }}{\sqrt{10} } \cdot \frac{ \sqrt{x^2}}{ \sqrt{x^{-7}}}=\] \[\LARGE \frac{\sqrt{45\cdot 2 \cdot 2 }}{\sqrt{10} } \cdot \frac{ x^{1/2} }{ x^{-\frac72 }} =\] \[\LARGE \frac{\sqrt{45\cdot 2^2 }}{\sqrt{10} } \cdot \left(x\right)^{\frac12-(-\frac72 )} =\] go on...

OpenStudy (anonymous):

Sqrt(x^3) is the same as x^(3/2) and Sqrt(x^-7) is equal to x^(-7/2). When dividing the x^(3/2) by x^(-7/2) we subtract the exponents 3/2 - (-7/2). Kreshnik you made an error it's the Sqrt(x^3) not of x^2 but the rest is ok.

OpenStudy (anonymous):

ahh sorry... @Kaden thanks for pointing that out :) ... @Courtney.soden just fix that and go on ;)

OpenStudy (anonymous):

lol and that wouldn't be 1/2 that would be 2/2 hahahah I'm so dumb lol

OpenStudy (anonymous):

No probs, I didn't even see the fact that I could post proper equations until after I posted.

OpenStudy (anonymous):

seems like she left. and Closed this. I'm out. Take care. :)

OpenStudy (anonymous):

You too, thanks for medal.